OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4

OpenStudy (anonymous):

Please Help!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@amistre64 @norasparkle4 @doggy @xxfaliottaxx @m@Kamille @itsloverboy @UmarsBack @undeadknight26 @elle_girl_0201 @PixieDust1 @Preetha

OpenStudy (giobott09):

The equation of the parabola: y = (1/12)(x - 6)² + 1

OpenStudy (giobott09):

Want me to show work?

OpenStudy (giobott09):

ocus of the parabola at (6, 4). horizontal line that is 6 units below the focus will be the directrix" directrix: y = -2 Determine the orientation of the parabola: The directrix is horizontal, so the parabola is vertical. The focus lies above the directrix, so the parabola opens upwards. Equation of up-opening parabola:  y = a(x - h)² + k The vertex is halfway between focus and directrix. vertex (6, 1) h = 6 k = 1 p = distance between focus and vertex = 3 a = 1/(4p) = 1/12

OpenStudy (anonymous):

Thanks ! :)

OpenStudy (giobott09):

@roche4321 Any Time :D

OpenStudy (anonymous):

Could you help me with another? ;o

OpenStudy (anonymous):

@Giobott09

OpenStudy (giobott09):

Ya

OpenStudy (mikurout):

@Giobott09 but the focus on the question is given at (0,-4).

OpenStudy (giobott09):

@mikurout yea but, 1/4 it's, so is negative, so we'll use -1/4 so let's plug all those values in vertex is at ( 2, -3/4) so distance from the vertex to the focus/directrix(x−2)2=4↑↓−14(y−−34)⟹(x−2)2=−1(y+34)↑vertex coordinates⟹(x−2)2−1=y+34⟹−(x−2)−34=y

OpenStudy (mikurout):

@Giobott09 why you will use it as -1/4. I think it's equation will be x^2 = -16y open down.

OpenStudy (anonymous):

Bit confused now..

OpenStudy (mikurout):

@roche4321 . I am right.

OpenStudy (giobott09):

@mikurout You can't be right when u haven't said the answer

OpenStudy (mikurout):

@Giobott09 The equation is x^2 = -16y

OpenStudy (anonymous):

@Giobott09 If the vertex is below the x axis would the parabola be shaped like a sad face?

OpenStudy (giobott09):

@mikurout Ask the smart one :p

OpenStudy (mikurout):

@Giobott09 Not required.

OpenStudy (anonymous):

Omg Someone please just say yes or no lol